# SimPro: A Simple Probabilistic Framework Towards Realistic Long-Tailed Semi-Supervised Learning

### v2

@@ 1 Introduction
Semi-supervised learning (SSL) offers a viable solution to the scarcity of labeled data by leveraging unlabeled data (Tarvainen & Valpola, 2017; Berthelot et al., 2019b; Miyato et al., 2018; Sohn et al., 2020). Common SSL algorithms typically generate pseudo-labels for unlabeled data to facilitate model training (Lee et al., 2013). However, real-world data often adheres to a long-tailed distribution, leading to a predominant focus on majority classes and resulting in imbalanced pseudo-labels (Li et al., 2022; Liu et al., 2019; Kang et al., 2020). This phenomenon, known as long-tailed semi-supervised learning (LTSSL), presents significant challenges in the field. Traditional LTSSL methods (Lai et al., 2022; Lee et al., 2021; Wei et al., 2022, 2021; Kim et al., 2020) assume consistency in class distributions between labeled and unlabeled data, an often unrealistic premise. In practice, class distributions can be inconsistent and unknown, especially as new data are continuously collected or from different tasks. This ongoing integration process can lead to significant shifts in class distributions.
@@
3. Comprehensive experiments on four commonly used benchmarks (CIFAR10-LT, CIFAR100-LT, STL10-LT, and ImageNet-127) and five distinct class distributions validate that our SimPro consistently achieves SOTA performance.
@@ 2 Related Work
In addressing Realistic LTSSL challenges, DASO (Oh et al., 2022) innovates by adapting the proportion of linear and semantic pseudo-labels to the unknown class distribution of unlabeled data. Its success largely depends on the discriminative quality of the representations, a factor that becomes less reliable in long-tailed distributions. ACR (Wei & Gan, 2023), on the other hand, attempts to refine consistency regularization by pre-defining distribution anchors and achieves promising results. CPE (Ma et al., 2024) trains multiple anchor experts where each is tasked to model one distribution. However, such anchor distribution-based approaches might not encompass all potential class distribution scenarios, and its complexity could hinder the broader application.
@@ 3.1 Preliminaries
In SSL, pseudo-labeling is a key technique for leveraging unlabeled data. This involves creating pseudo-labels for the unlabeled data using the model, followed by training model with both types of labels. This aligns with the Expectation-Maximization (EM) algorithm (Dempster et al., 1977), where the E-step generates pseudo-labels, and the M-step updates the parameters using the pseudo-labels, maximizing the likelihood function.
@@
where $\bm{\pi}$ and $\bm{\theta}$ are the parameters for the class and conditional distributions, respectively, and $\phi_{y}$ indicates the estimated class frequency in either the training or test set.
@@ 3.2 SimPro Framework
Here, we focus on estimating the parameters $\bm{\theta}$ and $\bm{\pi}$, pivotal for learning a discriminative model. Consequently, we concentrate on those terms dependent on $\bm{\theta}$ and $\bm{\pi}$, sidelining those independent of these parameters. The complete data log-likelihood is thus articulated as:
@@
For brevity, the superscript of the summation is omitted. The E-step involves generating soft pseudo-labels $P(y|x_{j};\bm{\theta}^{\prime},\bm{\pi}^{\prime})$ by employing the Bayes classifier (Prop. 2) under the current $\bm{\theta}^{\prime}$ and $\bm{\pi}^{\prime}$. In the subsequent M-step, these pseudo-labels are used alongside the one-hot labels of the labeled data to compute the cross-entropy loss.
@@
Prop. 3 illustrates that the regret bound is primarily governed by the first-order term of the estimation deviation. Additionally, it is inversely proportional to the ground truth $\bm{\phi}^{*}$, highlighting the learning challenges associated with imbalanced training data from a regret bound perspective.
@@
FixMatch (Sohn et al., 2020) 40.0±plus-or-minus\,$\pm$0.96 39.6±plus-or-minus\,$\pm$1.16 36.2±plus-or-minus\,$\pm$0.63 39.7±plus-or-minus\,$\pm$0.61 38.2±plus-or-minus\,$\pm$0.82 72.4±plus-or-minus\,$\pm$0.71 64.0±plus-or-minus\,$\pm$2.27
w/ CReST+ (Wei et al., 2021) 40.1±plus-or-minus\,$\pm$1.28 37.6±plus-or-minus\,$\pm$0.88 32.4±plus-or-minus\,$\pm$0.08 36.9±plus-or-minus\,$\pm$0.57 35.1±plus-or-minus\,$\pm$1.10 71.5±plus-or-minus\,$\pm$0.96 68.5±plus-or-minus\,$\pm$1.88
w/ DASO (Oh et al., 2022) 43.0±plus-or-minus\,$\pm$0.15 49.4±plus-or-minus\,$\pm$0.93 44.1±plus-or-minus\,$\pm$0.25 43.1±plus-or-minus\,$\pm$1.20 43.8±plus-or-minus\,$\pm$0.43 78.4±plus-or-minus\,$\pm$0.80 75.3±plus-or-minus\,$\pm$0.44
w/ ACR (Wei & Gan, 2023) 40.7±plus-or-minus\,$\pm$0.57 50.2±plus-or-minus\,$\pm$0.82 44.1±plus-or-minus\,$\pm$0.14 42.4±plus-or-minus\,$\pm$0.47 41.1±plus-or-minus\,$\pm$0.09 83.0±plus-or-minus\,$\pm$0.32 81.5±plus-or-minus\,$\pm$0.25
w/ SimPro 43.1±plus-or-minus\,$\pm$0.40 52.2±plus-or-minus\,$\pm$0.16 45.5±plus-or-minus\,$\pm$0.34 43.6±plus-or-minus\,$\pm$0.35 44.8±plus-or-minus\,$\pm$0.56 84.5±plus-or-minus\,$\pm$0.39 82.5±plus-or-minus\,$\pm$0.25
@@ 3.3 Implementation Details
where $\mathcal{L}_{l}$ and $\mathcal{L}_{u}$ represent the losses on labeled and unlabeled data, respectively. The hyper-parameter $\alpha$ acts as a scaling factor, the specifics of which will be elucidated later.
@@
For further clarity, the pseudo-code of SimPro is provided in Alg. 1. The modifications we made to the core training code, in comparison to FixMatch, are highlighted in bold. In the SimPro, we incorporate just a single additional line of code in the M-step to compute the closed-form solution of $\pi$ (Prop. 1). Furthermore, only four lines of code are modified to construct a Bayes classifier (Prop. 2) and to balance the loss between labeled and unlabeled data (denoted as $\alpha$). These minor yet crucial adjustments demonstrate that our SimPro framework is not only grounded in rigorous theoretical derivation but is also straightforward to implement in practice, exemplifying both simplicity and elegance.
@@ 4 Experiments
In this section, we first present the main results on various ReaLTSSL benchmarks in Sec. 4.1. More analysis, including the ablation studies and the visualization results, is presented in Sec. 4.2 to further evaluate the effectiveness of our SimPro. For detailed information regarding the experimental setup, please refer to App. A.
4.1 Results
Main results and comparison with SOTA baselines.
The results are presented in Tab. 1 (for CIFAR10-LT), Tab. 2 (for CIFAR100-LT and STL10-LT), and Tab. 3 (for ImageNet-127). It can be concluded that our method consistently outperforms the competitors across all distributions of unlabeled data and achieves SOTA performance. Notably, SimPro exhibits significant performance improvements on our two newly introduced distributions of unlabeled data: middle and head-tail. This substantiates the robust generalization capabilities of SimPro across various distributions that could potentially appear in real-world scenarios.
Moreover, we also conduct experiments on ImageNet-127, whose test dataset is imbalanced and consistent with the labeled data and unlabeled data. However, this is not suitable as a benchmark for long-tail learning, as biased classifiers tend to perform well in such scenarios, which is precisely what we aim to avoid. Therefore, we resample it to achieve a uniform test distribution ($\gamma_{t}=1$). The results highlight that our SimPro achieves substantial performance enhancements when evaluated against this balanced test dataset.
@@ 4.1 Results
Table 4: The impact of the predefined anchor distribution in ACR (Wei & Gan, 2023) and CPE (Ma et al., 2024) on CIFAR10-LT with $\gamma_{l}=150,N_{1}=500$, and $M_{1}=4000$. $\star$ denotes that we use the predefined anchor distributions to estimate $P(y|\bm{\pi})$ in our SimPro. See more analysis in the main text.
@@
4.2 Analysis
Ablation Study.
(a) Marginal distribution estimation. We first investigate the impact of the estimation for $P(y|\bm{\pi})$ (M-step, Prop. 1) and its usage in building the Bayes classifier for pseudo-label generation (E-step, Prop. 2). The results in Tab. 5 substantiate the high effectiveness and necessity of this estimation in driving the success of our approach, thereby validating our methodology both theoretically and practically.
Table 5: Abalation study of estimating the marginal distribution $P(y|\bm{\pi})$ in M-step (Eq. 7) and using it for constructing Bayes classifier in E-step (Eq. 5). We conduct the experiments on CIFAR10-LT with $\gamma_{l}=150,N_{1}=500$, and $M_{1}=4000$.
@@ 4.2 Analysis
Hyperparameter Sensitivity.
As outlined in App. A, we discover that reducing the threshold value improves performance for the CIFAR100-LT dataset. The rationale behind adjusting the confidence threshold is based on the observation that an increase in the number of classes typically results in a corresponding decrease in the confidence of the predictions. Consequently, it becomes necessary to lower the threshold to accommodate this change in confidence levels. A sensitivity analysis regarding the threshold value is presented in Fig. 4 of App. B. It is consistently observed across different settings that the optimal performance is achieved when the threshold is set at $t=0.2$ and $t=0.95$ for the CIFAR100-LT and CIFAR10-LT, respectively.
Visualization of Estimation Quality.
Our study includes a visualization of the estimated distribution quality in Fig. 5 of App. B. The vertical axis quantifies the Kullback-Leibler (KL) divergence, which measures the deviation of the estimated distribution from the ground truth. The results demonstrate that the accuracy of the estimation markedly enhances within just a few epochs of training. This empirical evidence validates the effectiveness of the theoretically derived estimation method of distribution, as outlined in Prop. 1.
5 Conclusion
We introduce SimPro, a novel probabilistic framework for realistic long-tailed semi-supervised learning (ReaLTSSL). This framework represents a pioneering advancement in the field by innovatively enhancing the Expectation-Maximization (EM) algorithm. The key innovation lies in the explicit separation of the estimation process for conditional and marginal distributions. In the M-step, this separation allows for the derivation of a closed-form solution for the marginal distribution parameters. Additionally, SimPro optimizes the parameters of the conditional distribution via gradient descent, facilitating the learning of a Bayes classifier. In the E-step. the Bayes classifier, in turn, generates high-quality pseudo-labels. SimPro is characterized by its elegant theoretical underpinnings and its ease of implementation, which requires only minimal modifications to existing codebases. Furthermore, we incorporate two innovative class distributions specifically for unlabeled data. These distributions provide a more comprehensive and realistic evaluation of ReaLTSSL algorithms. Empirical evidence from various benchmarks demonstrates that SimPro consistently delivers state-of-the-art performance, highlighting its robustness and superior generalization capabilities.
Appendix A Experimental Setup
Datasets.
To validate the effectiveness of SimPro, we employ four commonly used SSL datasets, CIFAR10/100 (Krizhevsky et al., 2009), STL10 (Coates et al., 2011), and ImageNet-127 (Fan et al., 2022). Following the methodology described in ACR (Wei & Gan, 2023), we denote the number of samples for each category in the labeled, unlabeled dataset as $N_{1}\!\geq\!\cdots\!\geq\!N_{K}$, $M_{1}\!\geq\!\cdots\!\geq\!M_{K}$, respectively, where $1,\cdots,K$ are the class indices. We define $\gamma_{l}$, $\gamma_{u}$, and $\gamma_{t}$ as the class imbalance ratios for labeled, unlabeled, and test data, respectively. We specify ‘LT’ for those imbalanced variants. These ratios are calculated as $\gamma_{l}\!=\!N_{1}/N_{K}$ and $\gamma_{u}\!=\!M_{1}/M_{K}$. The sample number of the $k$-th class follows an exponential distribution, expressed as $N_{k}\!=\!N_{1}\cdot\gamma_{l}^{-\frac{k-1}{K-1}}$ for labeled and $M_{k}\!=\!M_{1}\cdot\gamma_{u}^{-\frac{k-1}{K-1}}$ for unlabeled data. As illustrated in Fig. 1, for the CIFAR10/100 datasets, we constructed five class distributions to test the performance of different algorithms under more general settings. Regarding the STL10 dataset, due to the unknown ground-truth labels of the unlabeled data, we approach the experiments by controlling the imbalance ratio of the labeled data. For ImageNet-127, we follow the original setting in Fan et al. (2022) ($\gamma_{l}\!=\!\gamma_{u}\!=\!\gamma_{t}\approx 286$). Nevertheless, this approach does not serve as an appropriate benchmark for long-tail learning. In these scenarios, biased classifiers often exhibit high performance, which is exactly the outcome we seek to prevent. Consequently, we alse resample the test dataset to ensure a uniform distribution ($\gamma_{t}\!=\!1$). Following Fan et al. (2022), we reduce the image resolution to $32\times 32$ and $64\times 64$ in response to resource constraints.
Training hyper-parameters.
Our experimental setup mainly follows FixMatch (Sohn et al., 2020) and ACR (Wei & Gan, 2023). For example, we employ the Wide ResNet-28-2 (Zagoruyko & Komodakis, 2016) on CIFAR10/100 and STL10, and ResNet-50 (He et al., 2016) on ImageNet-127. All models are optimized with SGD. Several settings are slightly different from those in ACR: as outlined in Sec. 3.3, to achieve a balanced training, we set the ratio between unlabeled and labeled samples in each batch as $\mu\!=\!M/N$ ($8$ on CIFAR, $16$ on STL10-LT, and $2$ on ImageNet-127, ), where $M,N$ are the total sample numbers of unlabeled/labeled data. In contrast, $\mu$ is set as $2$ for all datasets in ACR. The effectiveness of this adjustment is validated in Tab. 6. The batch size for labeled data is $64$ on CIFAR10/100 and STL10-LT, and $32$ on ImageNet-127. To ensure a fair comparison, the training epochs are reduced to $86$ on CIFAR10/100 and STL10-LT, and $500$ on ImageNet-127. The initial learning rate $\eta$ is linearly scaled to $0.17$ on CIFAR10/100 and STL10-LT, and $0.01$ on ImageNet-127, which decays with a cosine schedule (Loshchilov & Hutter, 2017) as in ACR. Regarding the hyperparameter $\tau$ used in Eq. 14, we follow the guidelines from Menon et al. (2021) and set $\tau\!=\!2.0$ and $1.0$ for CIFAR10-LT/STL10-LT and CIFAR100-LT/ImageNet-127, respectively. For the confidence threshold $t$ in Eq. 1), we set $t\!=\!0.95$ on CIFAR10-LT/STL10-LT following Sohn et al. (2020). We adjust it to $0.2$ on CIFAR100-LT/Imagenet-127, as we observe that reducing the threshold enhances performance (Fig. 4).
Appendix B More Experimental Results

### v3

@@ 1 Introduction
Semi-supervised learning (SSL) offers a viable solution to the scarcity of labeled data by leveraging unlabeled data (Tarvainen & Valpola, 2017; Berthelot et al., 2019b; Miyato et al., 2018; Sohn et al., 2020). Common SSL algorithms typically generate pseudo-labels for unlabeled data to facilitate model training (Lee et al., 2013). However, real-world data often adheres to a long-tailed distribution, leading to a predominant focus on majority classes and resulting in imbalanced pseudo-labels (Liu et al., 2019; Kang et al., 2020; Du et al., 2024). This phenomenon, known as long-tailed semi-supervised learning (LTSSL), presents significant challenges in the field. Traditional LTSSL methods (Lai et al., 2022; Lee et al., 2021; Wei et al., 2022, 2021; Kim et al., 2020) assume consistency in class distributions between labeled and unlabeled data, an often unrealistic premise. In practice, class distributions can be inconsistent and unknown, especially as new data are continuously collected or from different tasks. This ongoing integration process can lead to significant shifts in class distributions.
@@
3. Comprehensive experiments on five commonly used benchmarks (CIFAR10/100-LT, STL10-LT, and ImageNet-127/1k) and five distinct class distributions validate that our SimPro consistently achieves SOTA performance.
@@ 2 Related Work
In addressing Realistic LTSSL challenges, DASO (Oh et al., 2022) innovates by adapting the proportion of linear and semantic pseudo-labels to the unknown class distribution of unlabeled data. Its success largely depends on the discriminative quality of the representations, a factor that becomes less reliable in long-tailed distributions. ACR (Wei & Gan, 2023), on the other hand, attempts to refine consistency regularization by pre-defining distribution anchors and achieves promising results. CPE (Ma et al., 2024) trains multiple anchor experts where each is tasked to model one distribution. However, such anchor distribution-based approaches might not encompass all potential class distribution scenarios, and their complexity could hinder the broader application.
@@ 3.1 Preliminaries
In SSL, pseudo-labeling is a key technique for leveraging unlabeled data. This involves creating pseudo-labels for the unlabeled data using the model and then training the model with both the pseudo-labeled and ground-truth data. This aligns with the Expectation-Maximization (EM) algorithm (Dempster et al., 1977), where the E-step generates pseudo-labels, and the M-step updates the parameters using the pseudo-labels, maximizing the likelihood function.
@@
where $\phi_{y}$ denotes the class frequency in the training or test set, $\bm{\pi}$ is the class distribution parameter and $\bm{\theta}$ is the parameter of $P(x|y)/P(x)$. Here we omit the parameter of $P(x)$ for simplicity. The detailed mathematical derivation is provided in App. A.
@@ 3.2 SimPro Framework
Here, we do not explicitly parameterize $P({x})$, as per the independence of parameters through conditional parameterization (Koller & Friedman, 2009). Thus, when ${x}$ is not a condition, the parameters of the relevant notions omit the parameters of $P({x})$, such as $P({x})$, $P({x}|y)$, $P({x},y)$, etc. According to Sec. 3.1, this may lead to a potential misunderstanding, as the equation $P({x})=\sum_{y}P({x}|y;\bm{\theta})P(y;\bm{\pi})$ seems to suggest that $P({x})$ is parameterized by $\bm{\theta}$ and $\bm{\pi}$, which is not the case. The detailed mathematical derivation is provided in App. A.
We focus on estimating the parameters $\bm{\theta}$ and $\bm{\pi}$, pivotal for learning a discriminative model. Consequently, we concentrate on those terms dependent on $\bm{\theta}$ and $\bm{\pi}$, sidelining those independent of these parameters.
The complete data log-likelihood is thus articulated as:
@@
The E-step involves generating soft pseudo-labels $P(y|x_{j};\bm{\theta}^{\prime},\bm{\pi}^{\prime})$ under the current $\bm{\theta}^{\prime}$ and $\bm{\pi}^{\prime}$. These soft pseudo-labels are specifically defined by Eq. 10, which is detailed in Prop. 2. In the subsequent M-step, these pseudo-labels are used alongside the one-hot labels of the labeled data to compute the cross-entropy loss.
@@
Prop. 3 illustrates that the regret bound is primarily governed by the first-order term of the estimation deviation. Additionally, it is inversely proportional to the ground truth $\bm{\phi}^{*}$, highlighting the learning challenges associated with imbalanced training data from a regret-bound perspective.
@@
FixMatch (Sohn et al., 2020) 40.0±plus-or-minus\,$\pm$0.96 39.6±plus-or-minus\,$\pm$1.16 36.2±plus-or-minus\,$\pm$0.63 39.7±plus-or-minus\,$\pm$0.61 38.2±plus-or-minus\,$\pm$0.82 FixMatch (Sohn et al., 2020) 72.4±plus-or-minus\,$\pm$0.71 64.0±plus-or-minus\,$\pm$2.27
w/ CReST+ (Wei et al., 2021) 40.1±plus-or-minus\,$\pm$1.28 37.6±plus-or-minus\,$\pm$0.88 32.4±plus-or-minus\,$\pm$0.08 36.9±plus-or-minus\,$\pm$0.57 35.1±plus-or-minus\,$\pm$1.10 w/ CReST+ (Wei et al., 2021) 71.5±plus-or-minus\,$\pm$0.96 68.5±plus-or-minus\,$\pm$1.88
w/ DASO (Oh et al., 2022) 43.0±plus-or-minus\,$\pm$0.15 49.4±plus-or-minus\,$\pm$0.93 44.1±plus-or-minus\,$\pm$0.25 43.1±plus-or-minus\,$\pm$1.20 43.8±plus-or-minus\,$\pm$0.43 w/ DASO (Oh et al., 2022) 78.4±plus-or-minus\,$\pm$0.80 75.3±plus-or-minus\,$\pm$0.44
w/ ACR (Wei & Gan, 2023) 40.7±plus-or-minus\,$\pm$0.57 50.2±plus-or-minus\,$\pm$0.82 44.1±plus-or-minus\,$\pm$0.14 42.4±plus-or-minus\,$\pm$0.47 41.1±plus-or-minus\,$\pm$0.09 w/ ACR (Wei & Gan, 2023) 83.0±plus-or-minus\,$\pm$0.32 81.5±plus-or-minus\,$\pm$0.25
w/ SimPro 43.1±plus-or-minus\,$\pm$0.40 52.2±plus-or-minus\,$\pm$0.16 45.5±plus-or-minus\,$\pm$0.34 43.6±plus-or-minus\,$\pm$0.35 44.8±plus-or-minus\,$\pm$0.56 w/ SimPro 84.5±plus-or-minus\,$\pm$0.39 82.5±plus-or-minus\,$\pm$0.25
@@ 3.3 Implementation Details
where $\mathcal{L}_{l}$ and $\mathcal{L}_{u}$ represent the losses on labeled and unlabeled data, respectively. The hyper-parameter $\alpha$ acts as a scaling factor, the specifics of which are elucidated later.
@@
For further clarity, the pseudo-code of SimPro is provided in Alg. 1. The modifications we made to the core training code, in comparison to FixMatch, are highlighted in bold. In the SimPro, we incorporate just a single additional line of code in the M-step to compute the closed-form solution of $\pi$ (Prop. 1). Furthermore, only four lines of code need to be modified to construct a Bayes classifier (Prop. 2) and to balance the loss between labeled and unlabeled data (denoted as $\alpha$). These minor yet crucial adjustments demonstrate that our SimPro framework is not only grounded in rigorous theoretical derivation but is also straightforward to implement in practice, exemplifying both simplicity and elegance.
@@ 4 Experiments
In this section, we first present the main results on various ReaLTSSL benchmarks in Sec. 4.1. More analysis, including the ablation studies and the visualization results, is presented in Sec. 4.2 to further evaluate the effectiveness of our SimPro. For detailed information regarding the experimental setup, please refer to App. B.
@@
ImageNet-1k $\gamma_{t}=1$
FixMatch (Sohn et al., 2020)
w/ ACR (Wei & Gan, 2023) 13.2 23.4
w/ ACR (Wei & Gan, 2023) 13.8 23.3
w/ SimPro 19.7 25.0
4.1 Results
Main results and comparison with SOTA baselines.
The results are presented in Tab. 1 (for CIFAR10-LT), Tab. 2 (for CIFAR100-LT and STL10-LT), and Tab. 3 (for ImageNet-127/1k). It can be concluded that our method consistently outperforms the competitors across all distributions of unlabeled data and achieves SOTA performance. Notably, SimPro exhibits significant performance improvements on our two newly introduced distributions of unlabeled data: middle and head-tail. This substantiates the robust generalization capabilities of SimPro across various distributions that could potentially appear in real-world scenarios.
Moreover, we also conduct experiments on ImageNet-127, whose test dataset is imbalanced and consistent with the labeled data and unlabeled data. However, this is not suitable as a benchmark for long-tail learning, as biased classifiers tend to perform well in such scenarios, which is precisely what we aim to avoid. Therefore, we resample it to achieve a uniform test distribution ($\gamma_{t}=1$). The results highlight that our SimPro achieves substantial performance enhancements when evaluated against this balanced test dataset. Beyond this, we further conduct experiments on ImageNet-1k to validate the performance of our method across a broader range of classes. The results in Tab. 3 demonstrate that our SimPro achieves state-of-the-art performance on ImageNet-1k.
Table 4: The impact of the predefined anchor distribution in ACR and CPE (Ma et al., 2024) on CIFAR10-LT with $\gamma_{l}=150,N_{1}=500$, and $M_{1}=4000$. $\star$ denotes that we use the predefined anchor distributions to estimate $P(y|\bm{\pi})$ in our SimPro. See more analysis in the main text and more results in App. C.
@@ 4.1 Results
Table 5: Ablation study of estimating the marginal distribution $P(y|\bm{\pi})$ in M-step (Eq. 7) and using it for constructing Bayes classifier in E-step (Eq. 5). We conduct the experiments on CIFAR10-LT with $\gamma_{l}=150,N_{1}=500$, and $M_{1}=4000$.